input output change in output
x y=2x-6
-2 2(-2)-6=-10 2
-1 ? -8=?
0 ? ? ? ?
1 ? ? ? ?
2 ? ? ? ?
question is complete this table?

Answer by drk(1908)

here is the completed table:
input . . . . . . . . output . . . . . . . . . . . . . . . . change in output
x . . . . . . . . . . . y = 2x - 6 . . . . . . . . . . . . . . . .y = ?
-2 . . . . . . . . . . 2(-2) - 6 =-10 . . . . . . . . . . . . . .
-1 . . . . . . . . . . .2(-1) - 6 = -8 . . . . . . . . . . .(+2)
0 . . . . . . . . . . . 2(0) - 6 = -6 . . . . . . . . . . . .(+2)
1 . . . . . . . . . . . .2(1) - 6 = -4. . . . . . . . . . . . (+2)
2 . . . . . . . . . . . .2(2) - 6 = -2. . . . . . . . . . . . (+2)